OpenStudy (anonymous):

what are the prime factors of 144

OpenStudy (anonymous):

144 = 12\(\times\)12 =(4\(\times\)3)\(\times\)(4\(\times\)3) = ((2\(\times\)2)\(\times\)3)\(\times\)((2\(\times\)2)\(\times\)3) = (2^4)(3^2)

OpenStudy (anonymous):

In other words:\[144=2^4*3^2 \]

OpenStudy (anonymous):

2 and 3 are both prime numbers.